Problem 1: Below are financial statements of FIN300.01 Inc: Balance Sheet: 12/31/17 Assets Cash and Marketable Securitics Accounts Reccivable Inventorics Total Current Asscts Net plant and equipment TOTAL ASSETS 2017 10 375 61.5 1,000 2016 80 31. 41.5 810 Liabilities and Equity Accounts Payable Notes Payable Accruals Total Current Liabilities Long Term Bonds TOTAL DEBT Preferred Stock Common Stock Retained carnings TOTAL COMMON EQUITY TOTAL LIABILITIES AND EQUITY 2017 140 110 310 2016 40 G0 130 230 1,064 40 130 700 830 130 766 896 2n 1,680 Income Statement: 12/31/17 Net Sales Operating Costs (excludes Dep/Amortization) EBITDA Depreciation Amortization Depreciation and Amortization EBIT Less Interest EBT Taxes (40%) NET INCOME (before Preferred Dividends) Preferred Dividends NET INCOME Common Dividends Addition to Retained Earnings 2017 2016 2,850 뿔뿔 100 90 90 263 100 312 187.2 1832 66.2 203 121.8 1178 53 64.8 Refer to FIN30001 Inc. what was the free cash flow in 2017 for FIN3 0001 Inc
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

. Home nert Page Layout Formulas Data Review View dd-Ins Cut aCopy Σ AutoSum Wrap Text в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Paste Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard AT194 Al Font Alignment Number Styles Cells Edting AK AL AM AN AO AP AQ AR AS AT AU AV AW 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 OCFEBITDEPRECIATION - TAXES OCF400 100 124.8 375.2 FCF FREE CASH FLOW OCF -NCS - NWC (NCS NET CAPITAL SPENDING AND NWC-NET WORKING CAPITAL) ANSWER FCF= 375.2 -230-110- 35.2 NOTE NCS-NET CAPITAL SPENDING NCS-CHANGE IN NET PLANT & EQUIPMENT + DEPRECIATION NCS = ( 1000-870) + 100 = 230 NWC = NET WORKING CAPITA-WORKING CAPITAL(2017)-WORKING CAPITAL (2016) NWC- 690-580 = 110 WORKING CAPITALCURRENT ASSETS CURRENT LIABILITIES 2017 = 2016 = 1000-310- 810-230- 690 580 | NpV 1 OCF CASH FLOW STAT N AR confidenceINDEX ../CURRENCY . LOAN CROSS OVER HEDGING MONEY MARKET , GM HPR AM TAX GROWTH SUST ( PREF AND DIV ρ 07:14 04-01-2019